The circular rug is placed on a square floor with side lengths of 14 feet. The rug touches all four walls. How much of the floor space is nit covered by the rug?
Answer by mananth(16946)

The circular rug is placed on a square floor with side lengths of 14 feet. The rug touches all four walls. How much of the floor space is nit covered by the rug?
The uncovered area = area of floor - area of rug
area of floor = s^2 = 14^2 = 196
The diameter of the rug is = the width of the floor
diameter = 14
radius = 7
Area of circle = pi * r^2
= 3.14 * 7^2
=153.86
196-153.86 =42.14 ft^2